Systems Staff Meeting Notes 3/8/99

Systems Staff Meeting Notes
March 8, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Ad Hoc
  1. Discussion on the phone setup - some departments have their phones set up so that incoming calls from the outside world have the option to directly dial an employee's extension and bypass any reception desk answering the calls. The general consensus in Systems is that this type of arrangement would not work for us, mainly due to the logging in of trouble calls. We may be able to crank up the number of rings allowed before the automated system answers for times when the phone is unattended.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. We have still not received a distribution list for the 20 Cows. Should be dispatched by this Wednesday.
  2. The InterDev programming books that Brian requested were ordered by Steve Carrico.
Liaisons
  1. Quiet.
Networking
  1. Outage this weekend - Sunday morning - they were making some changes on the CORE and left IPX out - which provides the connection from East to West. Bill had to come in - Cows1 was disabled with bogus password attempts - we can't set the number on individual account retries. The Libraries seem to catch these outages first - we must have the most active use of the network on campus. NERDC did not rebuild the CORE routers correctly, they're currently bridging to handle the problems.
  2. Wiring - Ashley has received two different quotes, Digital's was cheaper. They've quoted a rack for Systems at $1000, if we were to install it, we could do it for 10% of cost. Bill would like to have the final totals this week.
Software Projects Status
  1. ATC Standalone Catalog - Winston started working on it - downloaded 2 tables from FCLA - 1.7 Gig, which took 30 hours. He is now trying to make a backup. He's going to have to join a 2 million record table with a 3 million record table. Mark Hinnebush at FCLA may make some changes to speed up the download.
  2. LASSIE is coming along.
  3. IR Database tables have been settled, Ying is now working to carry out the plan.
  4. Winston & Brian were trying to install InterDev on Alena Aissing's student's machine to do Eastern European database - but it needs a memory upgrade. The upgrade has been ticketed to Everett.
  5. Library Order Tracking - an upcoming project which will be cloned out to LASSIE. Will track order and purchasing requests.
  6. Facilities Suite is on indefinite hold.
DP Coordinator
  1. Compustat tape status - Bobbie tried to map & copy the latest batch of ASCII cartridges, to no avail. She had Compustat send her documentation - which has still done no good. She even tried reading the EBCDIC cartridges, but they have so many header errors that they are unreadable.
Calendar
  1. Students are off all this week for Spring Break.
  2. Brian might be out the next two Thursdays.
  3. March 17 - St. Patrick's Day!


Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu